OpenStudy (luigi0210):

So just to clarify.. to fine the derivative of this:

OpenStudy (luigi0210):

\[\LARGE y=(\frac{x^2+4}{x^-4})^3\] It would be power rule, and the quotient rule for the inside?

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

ballz x 2

OpenStudy (lastdaywork):

Which is same as saying - first apply power rule ; then quotient rule.

OpenStudy (luigi0210):

Okay.. so this will be fun xD

OpenStudy (luigi0210):

\[\LARGE y=(\frac{x^2+4}{x^2-4})^3\] \[\LARGE \LARGE y'=3(\frac{x^2+4}{x^-4})^2*\frac{(x^2-4)(2x)-(2x)(x^2+4}{(x^2-4)^2}\]

OpenStudy (luigi0210):

And then simplification?

OpenStudy (shadowlegendx):

It looks so fun

OpenStudy (luigi0210):

That should be x^2-4 btw

OpenStudy (lastdaywork):

Yep..it's correct..

OpenStudy (luigi0210):

Thanks guys xD

OpenStudy (shadowlegendx):

I do no work but thanks :D

OpenStudy (shadowlegendx):

DID :P